State the domain and range for y = √(3x+4) – 5
Answer by stanbon(75887)

State the domain and range for y = √(3x+4) – 5
-----
Domain = ?
Solve 3x+4>=0
3x >= -4
x >= -4/3
Domain: All Real Numbers >= -4/3.
======================================
Range:
sqrt(3x+4) is always >= 0
So sqrt(3x+4)-5 >= -5
Range: All Real Numbers >= -5
==================================
Cheers,
Stan H.